OpenStudy (anonymous):

Find all solutions in the interval [0, 2π]: sin(x/2) = 1 - cosx

OpenStudy (anonymous):

What I have is: x=0, π, π/3, 5π/3--but I'm unsure about the latter two answers.

OpenStudy (anonymous):

x = 2 pi n x = 4 pi n + pi/3 x = 4 pi n + (5 pi)/3 for n in Z Check the attached image

OpenStudy (anonymous):

Haha, well using graphing technology or graphs isn't permitted but that does help verify my answers, I suppose. Thanks!

OpenStudy (anonymous):

hey hold it i got it, im gonna send u a pic with the equations and stuff

OpenStudy (anonymous):

Ah, cool, do our answers match up or am I wrong?

OpenStudy (anonymous):

I'm assuming the latter if you're sending me a picture of a sol'n, lol

OpenStudy (math&ing001):

You have cos(x)=1-2sin^2(x/2) sin(x/2) = 1 - cosx sin(x/2)=2sin^2(x/2) sin(x/2)(1-2sin(x/2)=0 sin(x/2)=0 or sin(x/2)=1/2

OpenStudy (anonymous):

Mhm, thanks.

OpenStudy (math&ing001):

yw